### v2

@@ Conclusions
This study demonstrates a novel approach to turbulence control by integrating explainable deep learning (XDL) with deep reinforcement learning (DRL) using the Shapley additive explanations (SHAP values). The SHAP values identify the most important coherent structures in the flow and they are capitalized on to discover control strategies for wall-bounded turbulence. By training DRL models in a minimal flow configuration with a single self-sustaining process, we successfully developed control strategies which can be deployed to much larger channels with multiple interacting SSPs.
Our analysis reveals several critical insights. While the DRL approach targeting Q events excels at reducing the classical Q events and streaks, it falls short in reducing those structures truly relevant for drag reduction. In contrast, the SHAP-based approach, relying on the SHAP values, yields a more effective control strategy, achieving $33.3\%$ drag reduction. This represents a $5\%$ improvement over the DRL approach directly targeting drag reduction, a $21.2\%$ improvement over Q-event-based control, and a $57\%$ improvement over the streak-based control. Furthermore, the SHAP-based control only requires half of the power input compared with the DRL approach targeting drag reduction.
@@
The framework introduced in this work is a promising entry point for the powerful and unique combination of DRL and XDL, where there is still a wide range of possible improvements such as: usage of a convolutional neural network in the DRL policy (instead of a multilayer perceptron), training using volumetric data instead of a single plane, or transfer learning from the small training domain to the large channel one. The use of XDL for DRL can have a great impact in a wide range of applications as it can be applied directly on experimental data [24], avoiding the ever growing requirements for computational power and high-resolution simulation. The SHAP reward introduced in this work provides new insights into the most critical mechanisms of turbulence and provides new paths to tame and harness the power of turbulent flows in many industrial (e.g. aerodynamics and food production), biological (e.g. active matter) and medical applications (e.g. personalised medicine).
@@ Methods
The flow field is entirely described by the velocity vector $\mathbf{u}(x,y,z,t)=(u,v,w)$. Direct numerical simulations (DNS) for training are carried out in a computational domain of dimensions $[L_{x},L_{y},L_{z}]=[2.67,1,0.8]$, i.e. the small channel configuration (SCC), and the evaluation of the resulting policies is performed in a box of size $[L_{x},L_{y},L_{z}]=[2\pi,1,\pi]$, i.e the large channel configuration (LCC). We consider Dirichlet boundary conditions at the lower wall and symmetry boundary conditions at the upper boundary:
@@
To perform our DNS we use the generic partial differential equation (PDE) solver Dedalus [31]. This solver uses a pseudo-spectral method, where solutions are expanded into Fourier series in the homogeneous ($x$ and $z$) directions and into Chebyshev polynomials into the inhomogeneous (wall-normal, $y$) direction. We expand solutions of the SCC into $[N_{x},N_{y},N_{z}]=[16,64,16]$ modes and we timestep the simulations using a third-order four-stage diagonally implicit Runge–Kutta and explicit Runge–Kutta (DIRK+ERK) scheme [32] with a constant time step $\Delta t^{+}\approx 0.039$ ($\Delta t=0.005$). These computational parameters are common in the literature for the moderate Reynolds number considered in this study [22]. The LCC simulations are performed expanding the solutions into $[N_{x},N_{y},N_{z}]=[64,64,32]$ modes using the same time-stepping scheme and time step as in the SCC case. These configurations (SCC and LCC) are used to perform long simulations ($t^{+}>20,000$) which form the dataset for training (and testing) required for the U-net models and the DRL framework.
@@
As briefly described in the introduction, a DRL set-up consists of two main elements that interact with each other: the environment, i.e. the system where we input actions, obtain rewards, and observe states, and the agent, i.e. the controller in charge of deciding the actions to take based on inputs from the environment. In the present work, as shown in Fig. 1, the environment corresponds to a numerical simulation performed using the Dedalus codebase [31], which accounts for the main computational cost in the DRL set-up. To address this, the Dedalus code is parallelized over several processors using the message-passing interface (MPI).
@@
with $\mathbb{E}$ denoting the expectation. Moreover, model-free agents can be classified as on-policy or off-policy. In on-policy methods, the optimization and policy update are based on the learning generated by the current agent. Off-policy algorithms use a replay buffer to store trajectories generated with previous policies, and these are used to perform the policy update. In the present work we have chosen the model-free, off-policy TD3 algorithm [33] as implemented in the Stable Baselines 3 library [34]. The combination of the Dedalus codebase with Stable Baselines 3 is particularly convenient since both of them are Python based. Note that this is a robust state-of-the-art DRL algorithm suitable for continuous control problems [33]. The reinforcement learning side of the set-up is much less computationally demanding and can be efficiently trained on a single CPU.
@@
In the physical configuration for the SCC described above, we defined the same number of controllers (agents) as grid points (modes) in the periodic directions, such that each agent sees only the state $(u,v)$ located at $y^{+}=15$ right above. The spatial average of the state is removed at each step, i.e. whenever information is communicated to the agent to avoid biasing the observations when evolving the flow between the controlled and uncontrolled cases (note that this is only required when information is passed to the agent). Consequently, each actuator is defined as a multi-agent reinforcement-learning pseudo-environment, resulting in $N_{x}\times N_{z}=256$ independent trajectories that the agent uses during the optimization step. The actions taken by the agents correspond to blowing and suction, i.e. positive/negative values of the vertical velocity at the wall, and they are limited to the range $[-u_{\tau},u_{\tau}]$. In order to ensure that no net mass is introduced by the controllers, we enforce a zero-mean action over all agents. This is done by removing the spatial average over all actuators before the control is applied in the DNS. The control interacts with the DNS every $\Delta t^{+}\approx 0.54$ ($\Delta t=0.07)$ and each episode consists of $3,000$ interactions, in agreement with prior studies [22]. Between consecutive agent interactions, the instantaneous values of the reward are calculated and stored at each DNS step. These stored values are then averaged to produce a single reward value.
@@
We train each model using 6 different initial conditions in the SCC. At the start of each episode a random number generator chooses which initial condition is used for that particular trajectory. Since the TD3 algorithm uses a deterministic policy in training, we introduce noise with amplitude $0.1u_{\tau}$ to favor exploration. During training in the SCC, relaminarization of the flow can be observed and in such cases we terminate the episodes once the flow is deemed to approach the laminar state sufficiently, i.e. once the drag reduction is over $55\%$. This is done to avoid providing the model with very large rewards for actions which no longer affect the flow on its path to relaminarization. The model is evaluated every 5 episodes on an additional initial condition not seen in training. The best performing model on the evaluated initial condition is then stored. Albeit several evaluation episodes could be considered, this process becomes expensive and a single evaluation initial condition proved to be sufficient to identify effective DRL models. Moreover, we perform several training runs to explore a wider range of control strategies. The Supplementary material shows the reward metrics during sample training runs for the different rewards considered. It can be observed that typically the model improves quickly at the start of the process and then discovers many similarly effective policies during training. Throughout the training, the model passes through parameter regions where better policies are found, but also where it unlearns. Each 100 episodes corresponds to roughly 96 hours of computing using four AMD Ryzen 9 7950X CPUs.
@@
The best policy for each of the four DRL rewards in the SCC is tested in 50 initial conditions randomly sampled from our database of $30,000$ instantaneous fields and unseen during training as reported in the Supplemetary material. The results show that all DRL policies outperform opposition control and for $t^{+}>500$ they achieve drag reductions of: $36.5\%$ for the DD reduction reward, $36.2\%$ for the Q-event-based reward, $34.4\%$ for the streak-based reward, and $37.4\%$ for the SHAP-based reward. It can also be observed that in this small channel, the SHAP-based reward produces the best performing policy in terms of reducing the Q events and streaks.
@@
For the first stage, a U-net architecture [35], $f_{u}$, is employed to predict the velocity fluctuation field in a future time step ($\mathbf{u}^{p}_{t+\Delta t}$) with $\Delta t^{+}=\Delta t\,u_{\tau}^{2}/\nu\approx 5$, from the true velocity field at the present time ($\mathbf{u}_{t}$): $\mathbf{u}^{p}_{t+1}=f_{u}\left(\textbf{u}_{t}\right)$. Both input and output fields have a size of $16\times 64\times 16$ grid points in the streamwise, wall-normal and spanwise directions, respectively. In addition, a periodic padding of $3$ is added in the streamwise and spanwise directions. The U-net has approximately $1.5$ million parameters, distributed on $4$ levels in the U-net, with $16$, $32$, $64$, and $128$ filters and a kernel of size $3\times 3$ from the first to the last level. Each level is composed of two blocks (3D convolutional layer+batch normalization+activation function) and one block in the decoder. The encoder levels are connected by average poolings with size $2$, while in the decoder, the levels are connected with transposed convolutions with $16$, $32$, and $64$ filters and kernel of size $3\times 3$ from top to bottom level. Each level of the encoder and the decoder are connected by concatenating the output of the encoder level with the output of the transposed convolution. The U-net is trained on a database of $30,000$ instantaneous flow fields, using $20\%$ for validation, until the prediction error is approximately $1\%$.
@@
Then, the importance of each grid point is calculated using additive-feature-attribution methods [25], which substitute the deep-learning model $f_{\rm MSE}$ by a surrogate linear model $g_{\rm MSE}$ defined as:
@@
where the importance, or Shapley additive explanations (SHAP) values, of each single grid point $i$ for the component $j$ of the velocity is defined by $\phi_{ji}$ and $N$ is the total number of grid points in a single snapshot. The parameter $z_{ji}$ is a boolean value which is 00 or $1$ in case of removing or including the information of the velocity component, $j$, in this specific grid point, $i$. The expected mean-squared error of the prediction when all the grid points are removed is defined by $\phi_{0}$. Note that the linear model presented in equation (14) is a local approximation of the MSE, and thus, it is updated for every snapshot. However, the computational cost of these values increases exponentially with the number of parameters, i.e. as $2^{N}$ [36]. To reduce the computational cost of the calculations, the prior knowledge of the mathematical definition of the architecture and the neurons can be exploited [37]. In the present work, the gradient-SHAP algorithm, an additive-feature-attribution method for differentiable models based on the expected-gradients method [38], is used to simplify the SHAP-value calculation as follows:

### v3

@@ Conclusions
This study demonstrates a novel approach to turbulence control by integrating explainable deep learning (XDL) with deep reinforcement learning (DRL) using the Shapley additive explanations (SHAP) values. The SHAP values identify the most important coherent structures in the flow and they are capitalized on to discover control strategies for wall-bounded turbulence. By training DRL models in a minimal flow configuration with a single self-sustaining process, we successfully developed control strategies which can be deployed accross Reynolds numbers and geometries through zero-shot generalization on a channel at $\textit{Re}_{\tau}=550$ and a zero-pressure-gradient turbulent boundary layer at $\textit{Re}_{\theta}=460$.
Our analysis reveals several critical insights. While the DRL approach targeting Q events excels at reducing the classical Q events and streaks, it falls short in reducing those structures truly relevant for drag reduction. In contrast, the SHAP-based approach, relying on the SHAP values, yields a more effective control strategy, achieving $33.7\%$ drag reduction. This represents a $5.6\%$ improvement over the DRL approach directly targeting drag reduction, a $25.3\%$ improvement over Q-event-based control, and a $61.2\%$ improvement over the streak-based control. Furthermore, the SHAP-based control only requires half of the power input compared with the DRL approach targeting drag reduction. This yields even more significant results, where the SHAP-based control achieves a net energy saving of $30.6\%$, which is $15.9\%$ better than controlling Q events, $18.1\%$ better than controlling the drag, $34.8\%$ better than opposition control, and $70\%$ better than controlling the streaks. Additionally, we show that the SHAP-based control is still the best performing policy in a turbulent channel flow at $\textit{Re}_{\tau}=550$ and for the ZPGTBL at $\textit{Re}_{\theta}=460$ without any retraining. The boundary layer case is particularly relevant, since the SHAP-based policy is the only control technique outperforming opposition control without any specific retraining.
@@
The framework introduced in this work is a promising entry point for the powerful and unique combination of DRL and XDL, where there is still a wide range of possible improvements such as: usage of a convolutional neural network in the DRL policy (instead of a multilayer perceptron), training using volumetric data instead of a single plane, or transfer learning from the small training domain to the large channel one. The use of XDL for DRL can have a great impact in a wide range of applications as it can be applied directly on experimental data [cremades2024identifying], avoiding the ever growing requirements for computational power and high-resolution simulation. The SHAP reward introduced in this work provides new insights into the most critical mechanisms of turbulence and provides new paths to tame and harness the power of turbulent flows in many industrial (e.g. aerodynamics and food production), biological (e.g. active matter) and medical applications (e.g. personalised medicine).
@@ Methods
The flow field is entirely described by the velocity vector $\mathbf{u}(x,y,z,t)=(u,v,w)$. Direct numerical simulations (DNS) for training are carried out in a computational domain of dimensions $L_{x},L_{y},L_{z}]=[2.67,1,0.8]$, i.e. the small channel configuration (SCC), and the evaluation of the resulting policies is performed in a box of size $L_{x},L_{y},L_{z}]=[2\pi,1,\pi]$, i.e the large channel configuration (LCC). We consider Dirichlet boundary conditions at the lower wall and symmetry boundary conditions at the upper boundary:
@@
To perform our DNS we use the generic partial differential equation (PDE) solver Dedalus [burns2020dedalus]. This solver uses a pseudo-spectral method, where solutions are expanded into Fourier series in the homogeneous ($x$ and $z$) directions and into Chebyshev polynomials into the inhomogeneous (wall-normal, $y$) direction. We expand solutions of the SCC into $N_{x},N_{y},N_{z}]=[16,64,16]$ modes and we timestep the simulations using a third-order four-stage diagonally implicit Runge–Kutta and explicit Runge–Kutta (DIRK+ERK) scheme [ascher1997stabilization] with a constant time step $\Delta t^{+}\approx 0.039$ ($\Delta t=0.005$). These computational parameters are common in the literature for the Reynolds number considered in this study [guastoni2023deep]. The LCC simulations are performed expanding the solutions into $N_{x},N_{y},N_{z}]=[64,64,32]$ modes using the same time-stepping scheme and time step as in the SCC case. These configurations (SCC and LCC) are used to perform long simulations ($t^{+}>20,000$) which form the dataset for training (and testing) required for the U-net models and the DRL framework. The set-up for the zero-shot generalization for a $\textit{Re}_{\tau}$=550 uses the same geometry as the LCC, and is expanded into $N_{x},N_{y},N_{z}]=[512,256,64]$ modes using the same timestepping scheme as the SCC with a constant time step $\Delta t=0.001$.
Turbulent boundary layer set-up
We consider a zero-pressure-gradient turbulent boundary layer (ZPGTBL) developing on a flat plate. In this configuration we use the same coordinates as in the SCC but $x$ is measured from the domain inflow, located at a distance $x_{0}$ from the leading edge. A local Reynolds number can be defined as $\textit{Re}_{\delta^{*}}=U_{\infty}\delta^{*}(x)/\nu$, where $U_{\infty}$ denotes the freestream velocity and $\delta^{*}(x)=\int_{0}^{L_{y}}(1-u/U_{\infty})dy$ is the displacement thickness. The equations are nondimensionalized by $U_{\infty}$ and $\delta^{*}_{0}$. We consider a computational domain of dimensions $[Lx,Lz,Ly]=[500,30,60]$ where $\textit{Re}_{\delta_{0}^{*}}=450$ is set at the inflow. We use the generic solver Dedalus [burns2020dedalus], which allows only for one inhomogeneous direction, and project our solution onto Fourier modes in the stream- and spanwise directions and Chebyshev polynomials in the wall-normal direction. We use $N_{x},N_{z},N_{y}]=[512,128,256]$ modes and to satisfy the periodicity in the $x$-direction despite the spatial development we use the fringe method detailed in [chevalier2007simson]. We consider a fringe of length $x_{\text{fringe}}=135$ at the end of the domain, with a rise region of $100$, a fall region of $30$ and a maximum amplitude of $0.8$. Turbulence is triggered by a random trip volume forcing applied at $x_{\text{trip}}=10$ from the start of the computational domain using the forcing detailed in [chevalier2007simson]. This results in a measured $Re_{\theta}=460$ at the start of the control region, where $\theta=\int_{0}^{L_{y}}u/U_{\infty}(1-u/U_{\infty})dy$
@@
Setting up a control environment requires the consideration of several elements, such as the position of the sensing plane, the control frequency and amplitude, as well as any parameters affecting the computation of the rewards. We set up our training configuration following the best practices in the literature. The distance from the wall of the sensing plane has been considered in [vinuesainfluence], concluding that $y^{+}=15$ results in the best performance, both for DRL for direct drag reduction and for opposition control. The frequency and intensity of the control are chosen in agreement with previous works [guastoni2023deep, Sonoda_Liu_Itoh_Hasegawa_2023]. The time delay used for the calculation of the SHAP values has been considered in [cremades2024identifying, cremades2024classically] concluding that $\Delta t^{+}=\{1,2,5,10\}$ do not significantly affect the identified coherent structures based on the SHAP values. More details about the calculation of the SHAP values can be found in the section below.
As briefly described in the introduction, a DRL set-up consists of two main elements that interact with each other: the environment, i.e. the system where we input actions, obtain rewards, and observe states, and the agent, i.e. the controller in charge of deciding the actions to take based on inputs from the environment. In the present work, as shown in Fig. 1, the environment corresponds to a numerical simulation performed using the Dedalus codebase [burns2020dedalus], which accounts for the main computational cost in the DRL set-up. To address this, the Dedalus code is parallelized over several processors using the message-passing interface (MPI).
@@
with $\mathbb{E}$ denoting the expectation. Moreover, model-free agents can be classified as on-policy or off-policy. In on-policy methods, the optimization and policy update are based on the learning generated by the current agent. Off-policy algorithms use a replay buffer to store trajectories generated with previous policies, and these are used to perform the policy update. In the present work we have chosen the model-free, off-policy TD3 algorithm [fujimoto2018addressing] as implemented in the Stable Baselines 3 library [stable-baselines3]. The combination of the Dedalus codebase with Stable Baselines 3 is particularly convenient since both of them are Python based. Note that this is a robust state-of-the-art DRL algorithm suitable for continuous control problems [fujimoto2018addressing]. The reinforcement learning side of the set-up is much less computationally demanding and can be efficiently trained on a single central-processing unit (CPU).
@@
In the physical configuration for the SCC described above, we defined the same number of controllers (agents) as grid points (modes) in the periodic directions, such that each agent sees only the state $(u,v)$ located at $y^{+}=15$ right above. The spatial average of the state is removed at each step, i.e. whenever information is communicated to the agent to avoid biasing the observations when evolving the flow between the controlled and uncontrolled cases (note that this is only required when information is passed to the agent). Consequently, each actuator is defined as a multi-agent reinforcement-learning pseudo-environment, resulting in $N_{x}\times N_{z}=256$ independent trajectories that the agent uses during the optimization step. The actions taken by the agents correspond to blowing and suction, i.e. positive/negative values of the vertical velocity at the wall, and they are limited to the range $u_{\tau},u_{\tau}]$. In order to ensure that no net mass is introduced by the controllers, we enforce a zero-mean action over all agents. This is done by removing the spatial average over all actuators before the control is applied in the DNS. The control interacts with the DNS every $\Delta t^{+}\approx 0.54$ ($\Delta t=0.07)$ and each episode consists of $3,000$ interactions, in agreement with prior studies [guastoni2023deep]. Between consecutive agent interactions, the instantaneous values of the reward are calculated and stored at each DNS step. These stored values are then averaged to produce a single reward value.
@@
We train each model using 6 different initial conditions in the SCC. At the start of each episode a random number generator chooses which initial condition is used for that particular trajectory. This initial condition is then evolved for 3000 interactions between the agent and the flow field. The actuation frequency is set to $\Delta t=0.07$ and the timestep of the simulation is set to $dt=0.005$. The reward is collected at each timestep and averaged before sharing it with the agent every 14 timesteps in the DNS. Moreover, the policy is updated every 30 interactions between the agent and the DNS. Since the TD3 algorithm uses a deterministic policy in training, we introduce noise with amplitude $0.1u_{\tau}$ to favor exploration. During training in the SCC, relaminarization of the flow can be observed and in such cases we terminate the episodes once the flow is deemed to approach the laminar state sufficiently, i.e. once the drag reduction is over $55\%$. This is done to avoid providing the model with very large rewards for actions which no longer affect the flow as it is on its path to relaminarization. The model is evaluated every 5 episodes on an additional initial condition not seen in training for evaluation. The best performing model on the evaluated initial condition is then stored. Although several evaluation episodes could be considered, this process becomes expensive and evaluating on a single initial condition proved to be sufficient to identify effective DRL models. Moreover, we perform several training runs to explore a wider range of control strategies. The Supplementary material shows the reward metrics during sample training runs for the different rewards considered. It can be observed that typically the model improves quickly at the start of the process and then discovers many similarly effective policies during training. Throughout the training, the model passes through parameter regions where better policies are found, but also where it unlearns. Each 100 episodes corresponds to roughly 96 hours of computing using four AMD Ryzen 9 7950X CPUs.
@@
The best policy for each of the four DRL rewards in the SCC is tested in 50 initial conditions randomly sampled from our database of $30,000$ instantaneous fields and unseen during training as reported in the Supplementary material. The results show that all DRL policies outperform opposition control and for $t^{+}>500$ they achieve drag reductions of: $36.5\%$ for the DD reduction reward, $36.2\%$ for the Q-event-based reward, $34.4\%$ for the streak-based reward, and $37.4\%$ for the SHAP-based reward. It can also be observed that in this small channel, the SHAP-based reward produces the best performing policy in terms of reducing the Q events and streaks.
The training in the small channel results in a deterministic policy $a_{t}=\mu(s_{t})$ where $a_{t}$ denotes the action and $s_{t}$ denotes the state. The deterministic policy in this case is a mapping $f:U\times V\rightarrow A$ where $U$ denotes all the possible values of $u^{\prime}(x,y^{+}=15)$ (fluctuations in the streamwise direction), $V$ denotes all the possible values of $v^{\prime}(x,y^{+}=15)$ (fluctuations in the wall-normal direction) and $A$ denotes all the possible values of the action. Since our policy is a non-linear mapping it can be evaluated in any system independently of the number of agents, i.e. each agent has access to the same policy and selects an action accordingly.
@@
For the first stage, a U-net architecture [ronneberger2015], $f_{u}$, is employed to predict the velocity fluctuation field in a future time step ($\mathbf{u}^{p}_{t+\Delta t}$) with $\Delta t^{+}=\Delta t\,u_{\tau}^{2}/\nu\approx 5$, from the true velocity field at the present time ($\mathbf{u}_{t}$): $\mathbf{u}^{p}_{t+1}=f_{u}\left(\textbf{u}_{t}\right)$. Both input and output fields have a size of $16\times 64\times 16$ grid points in the streamwise, wall-normal and spanwise directions, respectively. In addition, a periodic padding of $3$ is added in the streamwise and spanwise directions. The U-net has approximately $1.5$ million parameters, distributed on $4$ levels in the U-net, with $16$, $32$, $64$, and $128$ filters and a kernel of size $3\times 3$ from the first to the last level. Each level is composed of two blocks (3D convolutional layer+batch normalization+activation function) and one block in the decoder. The encoder levels are connected by average poolings with size $2$, while in the decoder, the levels are connected with transposed convolutions with $16$, $32$, and $64$ filters and kernel of size $3\times 3$ from top to bottom level. Each level of the encoder and the decoder are connected by concatenating the output of the encoder level with the output of the transposed convolution. The U-net is trained on a database of $30,000$ instantaneous flow fields, using $20\%$ for validation, until the prediction error is approximately $1\%$.
@@
Then, the importance of each grid point is calculated using additive-feature-attribution methods [lundberg2017unified], which substitute the deep-learning model $f_{\rm MSE}$ by a surrogate linear model $g_{\rm MSE}$ defined as:
@@
where the importance, or Shapley additive explanations (SHAP) values, of each single grid point $i$ for the component $j$ of the velocity is defined by $\phi_{ji}$ and $N$ is the total number of grid points in a single snapshot. The parameter $z_{ji}$ is a boolean value which is 0 or $1$ in case of removing or including the information of the velocity component, $j$, in this specific grid point, $i$. The expected mean-squared error of the prediction when all the grid points are removed is defined by $\phi_{0}$. Note that the linear model presented in equation (14) is a local approximation of the MSE, and thus, it is updated for every snapshot. However, the computational cost of these values increases exponentially with the number of parameters, i.e. as $2^{N}$ [jia2019]. To reduce the computational cost of the calculations, the prior knowledge of the mathematical definition of the architecture and the neurons can be exploited [cremades2025]. In the present work, the gradient-SHAP algorithm, an additive-feature-attribution method for differentiable models based on the expected-gradients method [erion2021], is used to simplify the SHAP-value calculation as follows:
@@
In the present setup, we use the aforementioned U-net without retraining to predict the SHAP values from full instantaneous velocity fields, even while control is applied. Although this might modify the SHAP values associated with the flow field under consideration, these values are computed with respect to an instantaneous spatial average. Therefore, suppressing SHAP values is expected to reduce fluctuations over the instantaneous mean flow, which on its own is unable to sustain turbulence. This further strengthens the robustness of the results: The SHAP values obtained for an uncontrolled turbulent flow yield good control strategies even when the underlying flow state is modified by a varying boundary condition at the wall.
While fryer2021shapley and huang2024failings raised significant concerns regarding the use of SHAP values for feature attribution, particularly their instability under correlation and tendency to assign importance to logically irrelevant features, these issues are mitigated in the present study due to the nature of turbulent flows. Dependencies in such physical systems are predominantly local, the risk of misleading global correlations is minimized. Furthermore, the inherent noise in feature attribution is suppressed by exploiting domain periodicity. By exploiting these physical constraints, we ensure that SHAP-based policies remain effective and provide a faithful representation of the underlying dynamics despite the theoretical approximation limits of the Shapley framework, as previously demonstrated for multiple causal systems in the supplementary material of cremades2024classically, where the SHAP values detected causal relationships in mediator, cofounder and collider systems. In fact, the SHAP values provide a sensitivity analysis for a deep-learning model. In general, this correlation does not imply causation. However, when the model is trained on time-resolved, high-quality data such as the temporal evolution of a turbulent flow the resulting SHAP analysis reflects the causal dynamics of the flow.